# Data-Driven Neural Estimation of Indirect Rate-Distortion Function

### v1

@@ I Introduction
Intuitively, since $S$ (and thus $d(S,y)$) is not observable, when performing coding, one can only work with some form of estimate of $d(S,y)$, and the conditional expectation $\bar{d}(x,y)$ turns out to be the choice. Further information-theoretic aspects of the indirect rate distortion problem, such as non-asymptotic and multi-terminal settings, have been studied in [11, 12, 13, 14, 15, 16] and references therein.
The information-theoretic fundamental trade-off of the indirect rate distortion problem is characterized by the indirect rate-distortion function (iRDF). As discussed in the preceding paragraph, the iRDF can be obtained as the RDF of an equivalent standard rate distortion problem. Analytical solutions of iRDF have been available for a few specific models only; see, e.g., [17, 1, 9, 2, 18]. Although for general models one can employ standard numerical algorithms such as the Blahut-Arimoto algorithm and its variants [19, 20, 21], in order to cope with modern applications where the statistical knowledge of the source and its observation may not be readily available, it is also of great interest to develop data-driven algorithms.
@@ III-B NEIRD Algorithm
For the problem in Section II, instead of directly optimizing the probability distribution $Q_{Y}$, we adhere to the idea of optimizing over a family of measurable mappings $T_{\theta}:\mathcal{Z}\to\mathcal{Y}$, which is further parameterized by a neural network with parameters $\theta\in\Theta$. This way, the probability distribution $Q_{Y}$ is the image measure induced by $T_{\theta}$, namely, $T_{\theta}\circ P_{Z}^{-1}$ where $P_{Z}$ is some basis probability distribution (such as a uniform distribution over a unit cube or a multivariate standard Gaussian distribution) and $\circ$ denotes the push-forward operator. The mapping approach has been applied to the evaluation of standard RDF in [26], and the existence of such measurable mapping is guaranteed by the Borel isomorphism theorem (see, e.g., [29]), from which we have that the measurable spaces $(\mathcal{Z},P_{Z})$ and $(\mathcal{Y},Q_{Y})$ are isomorphic to each other because they are both isomorphic to the $[0,1]$ interval with Lebesgue measure.
@@
Evaluating these expectations via high-dimensional integration is infeasible without exact statistical knowledge, so it is foreseeable that empirical expectation will be employed as approximate solution. Given i.i.d. datasets $(x_{i},s_{i})_{i=1}^{n}$ and $(z_{j})_{j=1}^{m}$ drawn from $P_{S,X}$ and $P_{Z}$, respectively, the expectations in (13) and (8) can be approximated as
@@ IV-A Synthetic Datasets
We consider two cases: jointly Gaussian model from [18] and binary classification model from [2].
@@ IV-B MNIST Dataset
Figure 3(a) displays the iRDF curve obtained via the NEIRD algorithm. Meanwhile, we also plot a theoretical solution claimed in [34], which is obtained under the extra assumption that $P_{S|X}$ is deterministic for the MNIST dataset, as $R(D)=\max\{H(S)-D,0\}$, a linear segment. [^2] We observe that the points generated by the NEIRD algorithm are clustered around end-points $(H(S),0)$ and $(0,H(S))$. This is due to the essentially constant-slope behavior of the iRDF for the MNIST, noting that the NEIRD algorithm computes the iRDF for each given slope $\lambda\geq 0$. In contrast, we emphasize that for our proposed NEIRD algorithm, there is no need to impose any assumption on the joint probability distribution $P_{S,X}$.

### v2

@@ I Introduction
Intuitively, since $S$ (and thus $d(S,y)$) is not observable, when performing coding, one can only work with some form of estimate of $d(S,y)$, and the conditional expectation $\bar{d}(x,y)$ turns out to be the choice. Further information-theoretic aspects of the indirect rate distortion problem, such as non-asymptotic and multi-terminal settings, have been studied in [11, 12, 13, 14, 15, 16] and references therein. The information-theoretic fundamental trade-off of the indirect rate distortion problem is characterized by the indirect rate-distortion function (iRDF). As discussed in the preceding paragraph, the iRDF can be obtained as the RDF of an equivalent standard rate distortion problem. Analytical solutions of iRDF have been available for a few specific models only; see, e.g., [17, 1, 9, 2, 18]. Although for general models one can employ standard numerical algorithms such as the Blahut-Arimoto algorithm and its variants [19, 20, 21], in order to cope with modern applications where the statistical knowledge of the source and its observation may not be readily available, it is also of great interest to develop data-driven algorithms.
@@ III-B NEIRD Algorithm
For the problem in Section II, instead of directly optimizing the distribution $Q_{Y}$, we adhere to the idea of optimizing over a family of measurable mappings $T_{\theta}:\mathcal{Z}\to\mathcal{Y}$, which is parameterized by a neural network with parameters $\theta\in\Theta$. For results on the complexity issue of the neural network $T_{\theta}$, see, e.g.,[29]. This way, the distribution $Q_{Y}$ is the image measure induced by $T_{\theta}$, namely, $T_{\theta}\circ P_{Z}^{-1}$ where $P_{Z}$ is some basis probability distribution (such as a uniform distribution over a unit cube or a multivariate standard Gaussian distribution) and $\circ$ denotes the push-forward operator. The mapping approach has been applied to the evaluation of standard RDF in [26], and the existence of such measurable mapping is guaranteed by the Borel isomorphism theorem (see, e.g., [30]), from which we have that the measurable spaces $(\mathcal{Z},P_{Z})$ and $(\mathcal{Y},Q_{Y})$ are isomorphic to each other because they are both isomorphic to the $[0,1]$ interval with Lebesgue measure.
@@
Evaluating these expectations via high-dimensional integration is infeasible without exact statistical knowledge, so it is foreseeable that empirical expectation will be employed as the approximate solution. Given i.i.d. datasets $(x_{i},s_{i})_{i=1}^{n}$ and $(z_{j})_{j=1}^{m}$ drawn from $P_{S,X}$ and $P_{Z}$, respectively, the expectations in (13) and (8) can be approximated as
@@ IV-A Synthetic Datasets
We consider two cases: jointly Gaussian model from [18] and binary classification model from [2]. For both synthetic datasets, the networks consists of five fully-connected layers.
@@ IV-B MNIST Dataset
In this case, $g_{\hat{\theta}}$ consists of three convolution layers, and a generative adversarial network is used to build $T_{\theta}$. Figure 3(a) displays the iRDF curve obtained via the NEIRD algorithm. Meanwhile, we also plot a theoretical solution claimed in [35], which is obtained under the extra assumption that $P_{S|X}$ is deterministic for the MNIST dataset, as $R(D)=\max\{H(S)-D,0\}$, a linear segment. [^2] We observe that the points generated by the NEIRD algorithm are clustered around end-points $(H(S),0)$ and $(0,H(S))$. This is due to the essentially constant-slope behavior of the iRDF for the MNIST, noting that the NEIRD algorithm computes the iRDF for each given slope $\lambda\geq 0$. In contrast, we emphasize that for our proposed NEIRD algorithm, there is no need to impose any assumption on the joint probability distribution $P_{S,X}$.